### v1

# Infinite systems of equations in abelian and nilpotent groups111 This work was supported by the Russian Science Foundation, project no. 22-11-00075

Abstract
Every abelian (and even every nilpotent) group contains a solution of any finite unimodular system of equations over itself. However, this is not true for infinite systems. We deduced a criterion for a periodic abelian group to contain a solution of any infinite unimodular system of equations over itself. Using this criterion, we show that nilpotent groups of bounded period also contain solutions of all infinite unimodular systems of equations over themselves. Solvability of every nonsingular infinite system of equations in each divisible nilpotent group is shown as well.
1 Introduction
In this work $\mathbb{Z}_{p}$ for a prime number $p$ denotes the field of order $p$.
The present article is devoted to equations over groups.
Definition.
Let $G$ be a group. An equation in variables $x_{1},\ldots,x_{n}$ over $G$ is an expression (possibly with coefficients from $G$) $w(x_{1},\ldots,x_{n})=1$, where $w$ is an element of the free product $G*F(x_{1},\ldots,x_{n})$, in which $F(x_{1},\ldots,x_{n})$ is the free group with basis $x_{1},\ldots,x_{n}$.
The equation $w(x_{1},\ldots,x_{n})=1$ is solvable in the group $\widetilde{G}$ if $\widetilde{G}\supset G$ and $\widetilde{G}$ contains a solution of this equation (i.e. there are elements $\widetilde{g}_{1},\ldots,\widetilde{g}_{n}\in\widetilde{G}$ such that $w(\widetilde{g}_{1},\ldots,\widetilde{g}_{n})=1$). The group $\widetilde{G}$ is called a solution group. Equivalently, $w=1$ is solvable in $\widetilde{G}$ if there is a homomorphism $G*F(x_{1},\ldots,x_{n})\to\widetilde{G}$ which is injective on $G$ and sends $w$ to $1$.
If $w=1$ is solvable in some group $\widetilde{G}$, we say that the equation $w=1$ over $G$ is solvable.
The solvability of a (finite or an infinite) system of equations (possibly in an infinite set of variables) over a group is defined likewise.
The study of equations over groups has a long history but continues to this day. For example, see [3, 4, 14, 17, 23, 27, 1, 2, 6, 7, 8, 10, 11, 12, 13, 15, 16, 18, 9, 21, 24, 25, 28, 29]. There is a survey on equations over groups as well, see [26].
Here we are interested in nonsingular systems of equations.
Definition.
Let $\{w_{1}=1,\ldots,w_{k}=1\}$ be a finite system of equations in variables $\{x_{1},\ldots,x_{n}\}$ over a group $G$. Take the integer matrix $(a_{ij})$ of size $k\times n$, where $a_{ij}$ is the exponent sum of the variable $x_{j}$ in the word $w_{i}$ (we also call this number an exponent sum of the variable $x_{j}$ in the equation $w_{i}=1$). For example, the exponent sum of $x$ in the equation $x^{-1}gyx^{2}g^{2}x^{-3}y^{2}=1$ equals $-2$.
The system $\{w_{1}=1,\ldots,w_{k}=1\}$ is called nonsingular if the rows of the matrix $(a_{ij})$ are linearly independent over $\mathbb{Q}$. Equivalently, if these rows are independent as elements of a $\mathbb{Z}$-module, this is if no combination of these rows is equal to zero, apart from the combination where every coefficient is zero.
Let $p$ be a prime number. Let $\widetilde{a}_{ij}$ be a residue class of $a_{ij}$ modulo $p$. The system $\{w_{1}=1,\ldots,w_{k}=1\}$ is called $p$-nonsingular if the rows of the matrix $(\widetilde{a}_{ij})$ are linearly independent over $\mathbb{Z}_{p}$.
Let $\pi$ be some set of prime numbers (maybe not of all prime numbers). A finite system of equations is called $\pi$-nonsingular if it is $p$-nonsingular for every prime number $p\in\pi$. By definition $\varnothing$-nonsingular system is just a nonsingular system.
A finite system of equations is called unimodular if it is $\Pi$-nonsingular, where $\Pi$ is a set of all prime numbers.
An infinite system of equations is called nonsingular ($p$-nonsingular, $\pi$-nonsingular respectively), if its every finite subsystem is nonsingular ($p$-nonsingular, $\pi$-nonsingular respectively).
In particular, a single equation $w(x)=1$ is a single variable is nonsingular ($p$-nonsingular, unimodular respectively) if the exponent sum of $x$ in $w=1$ is not equal to zero (is not divisible by $p$, is equal to ±$\pm 1$ respectively).
One can consider an integer matrix of exponent sums of variable for an infinite system $\{w_{j}=1\}_{j\in J}$ of equations in variables $\{x_{i}\}_{i\in I}$ over a group $G$ as well. It is a matrix of size $I\times J$ with finitely supported rows, i.e. rows which have only a finite number of non-zero elements. Clearly, a system of equations is nonsingular if and only if the rows of the corresponding matrix are linearly independent over $\mathbb{Q}$. Analogously, the system is $p$-nonsingular if and only if the rows containing the corresponding residue classes of the elements of the matrix are linearly independent over $\mathbb{Z}_{p}$.
If a system of equations over a group is $p$-nonsingular for some prime number $p$ then it is also nonsingular, as having a nontrivial combination of integer rows over $\mathbb{Z}$, which is equal to zero, implies that there is a likewise combination in which at least one coefficient is not divisible by $p$. Particularly, unimodular systems of equations are nonsingular. Nonsingularity of a finite system of equations can be interpreted as existence of a non-zero minor of rank $k$ in the matrix of exponent sums of variables, where $k$ is the number of equations. Similarly, $p$-nonsingularity can be seen as existence of a likewise minor which is not divisible by $p$. Hence, if a system of equations is finite, then its nonsingularity implies its $p$-nonsingularity for some prime number $p$. This, however, is not true for infinite systems: the system $\{x_{p}^{p}=1\},$ where $p$ ranges over all prime numbers, is nonsingular but is $p$-singular for every prime number $p$.
Any nonsingular system of equations is solvable over following types of groups:
  • finite groups [10];
  • locally residually finite groups (follows from the previous item and some simple arguments);
  • hyperlinear groups (it is yet unknown if any group is such) [25];
  • locally indicable groups [11], i.e. groups whose every nontrivial finitely generated subgroup admits an epimorphism onto the additive group $\mathbb{Z}$;
  • $p$-nonsingular systems are solvable over locally $p$-indicable groups [9], see also [19]. Locally $p$-indicable group is a group whose every nontrivial finitely generated subgroup admits an epimorphism onto the additive group of $\mathbb{Z}_{p}$;
  • there is a conjecture (yet neither proved nor disproved) that all nonsingular systems of equations are solvable over every group [11].
In fact, the solvability of a system $\{w_{i}=1\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over a group $G$ is equivalent to the injectivity of a natural homomorphism
$G\to\left(G\ast F(X)\right)/\langle\langle\{w_{i}\}_{i\in I}\rangle\rangle,$
where $X=\{x_{j}\}_{j\in J}$. This causes, in particular, the locality property: the system of equations over a group is solvable if and only if its every finite subsystem is solvable. That is why it suffices to consider only finite systems to research solvability of nonsingular equations over groups.
Consider a finite $\pi$-nonsingular system of equations over a group and the matrix of the exponent sums of variables in these equations. Suppose that there are $k$ equations and $n$ variables in the system and $k<n$ (i.e. the matrix of the system is not square). Using transformations (i.e. multiplying one equation by another or by an inverse of another and rearranging equations) of equations in the system (which lead to transformations of rows in the matrix) and changes of variables (which lead to transformations of columns in the matrix), we can transform the matrix to a triangular matrix in which the first $k$ columns modulo $p$ are linearly independent over $\mathbb{Z}_{p}$ for each $p\in\pi$. Therefore this transformed system of equations remains $\pi$-nonsingular even after elimination of all variables except the first $k$ ones (for example, the other variables can be changed to $1$). Thus, any finite $\pi$-nonsingular system of equations can be reduced to a likewise system in which there are as many equations as there are variables (i.e. the matrix of the system is square). So it can be assumed for convenience that finite nonsingular, $\pi$-nonsingular or unimodular systems are ‘‘square’’.
One can also research solving systems of equations in groups which are similar to the initial group. For example, one can try to solve a system of equations over an abelian group in an abelian group as well (where the best case is being able to solve the system in the initial group itself). Let us start with finite systems.
If a group is finite, then every finite nonsingular system of equations over this group is solvable in a finite group as well [10]. If a group is abelian, then every finite nonsingular system of equations over this group is solvable in an abelian group, namely in the divisible hull of the group. For example, one can solve the system by Cramer’s rule (having in mind that it is possible to divide in a divisible abelian group, not necessarily uniquely). Moreover, if a finite system of equations over an abelian groups is unimodular, then the system is solvable in this abelian group itself. This is also true for all nilpotent group: it is a special case of Shmel’kin’s theorem. Before we formulate this theorem, let us remind a couple of definitions.
Definition.
Let $\pi$ be a set of some prime numbers. A group $G$ is called $\pi$-divisible if any element of $G$ is a $p$th power of another element of $G$ for each prime $p\in\pi$.
Particularly, a $\varnothing$-divisible group is just any group.
Definition.
Let $\pi$ be a set of some prime numbers. A group $G$ is called $\pi$-torsion free if it has no elements whose order is contained in $\pi$.
Particularly, a $\varnothing$-torsion-free group is just any group.
Let us denote the complement of $\pi$ in the set of all prime numbers by $\pi^{\prime}$.
Shmel’kin’s theorem ([27]).
Let $\pi$ be a set of some prime numbers. Suppose that $G$ is a $\pi$-divisible $\pi$-torsion-free locally nilpotent group. Then every square finite $\pi^{\prime}$-nonsingular system of equations over $G$ has a unique solution in $G$.
It follows that non-square finite $\pi^{\prime}$-nonsingular systems have a solution (although not unique) in $G$ as well, as discussed above.
The case where a system of equations over a group does not have solutions in groups from the same class, is also possible. For instance, there is a unimodular equation over a metabelian group such that the equation is unsolvable in metabelian groups [16].
The solution group of an infinite system of equations can differ from solution groups of finite systems. For example, an infinite nonsingular system of equations over a finite group can be unsolvable in finite groups, as shown by the system $\{x_{1}^{2}=a,x_{2}^{2}=x_{1},x_{3}^{2}=x_{2},\ldots\}$ over $\langle a\rangle_{2}$, i.e. cyclic group of order $2$. That is why for finding solutions of infinite systems of equations in groups with ceratin properties it is not enough to consider only finite subsystems.
The articles [16, 23] are concerned with finding solutions of infinite systems of equations over solvable groups in solvable groups as well.
Theorem ([16]).
Suppose that a group $G$ has a subnormal series with abelian factors
$G=G_{1}\triangleright G_{2}\triangleright\ldots\triangleright G_{n}\triangleright G_{n+1}=\{1\},$
in which all factors except the last one are torsion free. Then any (including the infinite ones) nonsingular system of equations over $G$ has a solution in some group $\widetilde{G}$, which also has a subnormal series with abelian factors
$\widetilde{G}=\widetilde{G}_{1}\triangleright\widetilde{G}_{2}\triangleright\ldots\triangleright\widetilde{G}_{n}\triangleright\widetilde{G}_{n+1}=\{1\},$
in which all factors except the last one are torsion free. If $G_{n}$ is $\pi$-torsion free, then $\widetilde{G}$ can be chosen such that $\widetilde{G}_{n}$ is $\pi$-torsion free as well.
Theorem ([23]).
Suppose that a group $G$ has a subnormal series with abelian factors
$G=G_{1}\triangleright G_{2}\triangleright\ldots\triangleright G_{n}\triangleright G_{n+1}=\{1\},$
in which all factors except the last one are torsion free. Then $G$ embeds into a group $\widehat{G}$, which:
  • has a subnormal series with abelian factors
    $\widehat{G}=\widehat{G}_{1}\triangleright\widehat{G}_{2}\triangleright\ldots\triangleright\widehat{G}_{n}\triangleright\widehat{G}_{n+1}=\{1\},$
    in which all factors except the last one are torsion free;
  • contains a solution of any nonsingular system of equations over itself. Particularly, $\widehat{G}$ contains a solution of every such system over $G$.
If $G_{n}$ is $\pi$-torsion free, then $\widehat{G}_{n}$ is $\pi$-torsion free as well.
Theorem ([23]).
Suppose that a group $G$ has a subnormal series with abelian factors
$G=G_{1}\triangleright G_{2}\triangleright\ldots\triangleright G_{n}\triangleright G_{n+1}=\{1\},$
in which all factors except the last one are $p^{\prime}$-torsion free. Then $G$ embeds into a group $\widehat{G}$, which:
  • has a subnormal series with abelian factors
    $\widehat{G}=\widehat{G}_{1}\triangleright\widehat{G}_{2}\triangleright\ldots\triangleright\widehat{G}_{n}\triangleright\widehat{G}_{n+1}=\{1\},$
    in which all factors except the last one are $p^{\prime}$-torsion free;
  • contains a solution of any $p$-nonsingular system of equations over itself. Particularly, $\widehat{G}$ contains a solution of every such system over $G$.
If $G_{n}$ is $\pi$-torsion free, then $\widehat{G}_{n}$ is $\pi$-torsion free as well.
The present article studies solvability of infinite systems of equations in abelian and nilpotent groups: it is researched which abelian and nilpotent groups contain solutions of all infinite unimodular systems of equations over themselves (having in mind that every nilpotent group has a solution of every finite unimodular system of equations over itself). It turns out that not even all abelian groups have this property. For instance, the group $\mathbb{Z}_{2}\oplus\mathbb{Z}_{3}\oplus\mathbb{Z}_{5}\oplus\ldots$ has no solutions of the system of equations (in the additive notation)
$\begin{cases}x+2y_{2}=(1,0,0,\ldots)\\ x+3y_{3}=(0,1,0,\ldots)\\ x+5y_{5}=(0,0,1,\ldots)\\ \ldots\end{cases}$
In Section 2, we prove the following criterion.
Theorem 1.
Suppose that $A$ is a periodic abelian group. Then every unimodular system of equations over $A$ is solvable in $A$ if and only if the reduced part of $A$ has bounded period.
In Section 3, using this criterion and the methods of [27], we prove the following result, which is a partial generalization of Shmel’kin’s theorem to the infinite systems of equations for nilpotent groups of bounded period.
Theorem 2.
Suppose that $G$ is a nilpotent group of bounded period. Then every (not even necessarily finite) unimodular system of equations over $G$ is solvable in $G$ itself.
Also, in this section we prove the same generalization of Shmel’kin’s theorem for divisible nilpotent group (not even necessarily torsion free).
Theorem 3.
Suppose that $G$ is a divisible nilpotent group. Then every (not even necessarily finite) nonsingular system of equations over $G$ is solvable in $G$ itself.
In Section 4, we formulate some open questions on solvability of infinite systems of equations in groups.
The author thanks the Theoretical Physics and Mathematics Advancement Foundation ‘‘BASIS’’. The author also thanks Anton Klyachko for valuable remarks and the mentorship.
2 Abelian groups
In this section, we look for solutions of systems of equations over an abelian group in an abelian group. Hence, if we rearrange some coefficients and variables in equations of a system, we get an equivalent (for our purpose) system. So, we can assume that variables commute with each other and with coefficients. In this case we can use the additive notation instead of the multiplicative one. Thus, in this section we may assume that each equation in a system has the form $k_{1}x_{1}+\ldots+k_{n}x_{n}=a$ for some $n\in\mathbb{N}$ (not necessarily the same $n$ for all equations), where $a$ is the coefficient of the equation, $x_{i}$ are variables and $k_{i}$ are integer numbers. The other systems of equations are equivalent to these (in the sense of searching for solutions in abelian groups). In this case $k_{i}$ is the exponent sum of $x_{i}$ in the equation. In other words, the left hand sides of such equations are the rows of exponent sums of variables.
Each abelian groups is a direct sum of a divisible abelian group and a reduced abelian group. Any nonsingular system of equations over a divisible abelian group $A$ has a solution in $A$ itself, as the following lemma shows.
Lemma 1.
Suppose that $A$ is a divisible abelian group. Then every (including the infinite ones) nonsingular system of equations over $A$ is solvable in $A$ itself.
Proof.
Let $\{m_{i}=a_{i}\}_{i\in I}$ be a nonsingular system of equations over $A$, where $a_{i}\in A$, while $m_{i}$ is an element of the $\mathbb{Z}$-module $\sum_{j\in J}\mathbb{Z}\cdot x_{j}$, which can be seen as an integer row.
The natural mapping $A\to\left(A\oplus\sum\mathbb{Z}\cdot x_{j}\right)/\langle\{m_{i}-a_{i}\}\rangle$ is an embedding because the rows $m_{i}$ are independent as elements of $\sum\mathbb{Z}\cdot x_{j}$. As $A$ is a divisible abelian group, it is a direct summand of $\left(A\oplus\sum\mathbb{Z}\cdot x_{j}\right)/\langle\{m_{i}-a_{i}\}\rangle$. The images of the elements $x_{j}$ under the projection onto $A$ form the solution of $\{m_{i}=a_{i}\}$ in $A$, as needed. ∎
That is why from now we consider only reduced abelian groups.
2.1 Negative case
Let us start with reduced abelian groups which have no solutions of some unimodular systems of equations over themselves.
Proposition 1.
Suppose that $A$ is a reduced abelian $p$-group. Suppose that the period of $A$ is not bounded. Then there is a unimodular system of equations over $A$, which has no solutions in $A$.
Proof.
Let us recall that the height of an element $a$ in an abelian $p$-group is the maximal number $k$, such that the equation $p^{k}x=a$ is solvable in this group, or infinity if there is no maximal number among such numbers $k$.
Note that $A$ can be assumed to have no elements of infinite height aside from 00. Indeed, if $A$ contains a solution of every unimodular system of equations over itself, then this fact is also true for any quotient group of $A$. In particular, it is true for the first Ulm factor $\bar{A}^{0}$ of $A$, i.e. quotient group of $A$ by the group $A^{1}$ consisting of elements of $A$ which have infinite height.
If $A$ has no elements of infinite height (except 00), then $A^{1}=\{0\}$ and $A\cong\bar{A}^{0}$. And if $A^{1}\neq\{0\}$, then note that in this case $\bar{A}^{0}$ has unbounded period as well and has no elements of infinite height aside from 00 [20, §27].
Therefore, we can assume that $A$ is an abelian $p$-group of unbounded period and without non-zero elements of infinite height (notice that the last property is stronger than the property of being reduced). Now we show that for every natural number $N$ the group $A$ has a cyclic direct summand of order not less than $p^{N}$.
Indeed, take an arbitrary element $a\in A$ of order $p^{N}$. The group $A$ is separable, as it has no elements of infinite height [5, §65], hence $a$ lies in a direct summand of $A$ which is a direct sum of some cyclic groups. Since the order of $a$ is $p^{N}$, at least one of these cyclic groups has order not less than $p^{N}$.
$A$ has a cyclic direct summand $\langle a_{1}\rangle_{p^{k_{1}}}$ of order not less than $p$:
$A=\langle a_{1}\rangle_{p^{k_{1}}}\oplus B_{1}.$
Now note that $B_{1}$ is also an abelian $p$-group of unbounded period without non-zero elements of infinite height. It has a cyclic direct summand of order not less than $p^{2k_{1}+1}$. We get the following decomposition of $A$:
$A=\langle a_{1}\rangle_{p^{k_{1}}}\oplus\langle a_{2}\rangle_{p^{k_{2}}}\oplus B_{2},$
where $k_{2}>2k_{1}$. Similarly, $B_{2}$ has a direct summand $\langle a_{3}\rangle_{p^{k_{3}}}$ such that $k_{3}>2k_{2}$ and so on.
On the $n$th step we get the following decomposition of $A$:
$A=\langle a_{1}\rangle_{p^{k_{1}}}\oplus\ldots\oplus\langle a_{n}\rangle_{p^{k_{n}}}\oplus B_{n},$
where $k_{i+1}>2k_{i}$ for each $i$ from $1$ to $n$.
Consider the following system of equations over $A$:
$\{x_{i}-p^{k_{i}-k_{i-1}}x_{i+1}=a_{i}\mid i\in\mathbb{N}\},$
where $k_{0}=0$. This system is unimodular, as can be seen from the matrix of exponent sums of the variables in the equations:
$\begin{pmatrix}1&-p^{k_{1}}&0&\dots\\ 0&1&-p^{k_{2}-k_{1}}&\dots\\ 0&0&1&\dots\\ \dots&\dots&\dots&\dots\\ \end{pmatrix}$
Suppose that this system has a solution $\{\widetilde{x}_{i}\}$. Then the following equalities hold:
$\widetilde{x}_{1}-p^{k_{i}}\widetilde{x}_{i+1}=a_{1}+p^{k_{1}}a_{2}+\ldots+p^{k_{i-1}}a_{i}.$
Take a look at the $i$th of these equations in the following decomposition of $A$:
$A=\langle a_{1}\rangle_{p^{k_{1}}}\oplus\ldots\oplus\langle a_{i}\rangle_{p^{k_{i}}}\oplus B_{i}.$
In the component $\langle a_{i}\rangle_{p^{k_{i}}}$ the element $p^{k_{i}}\widetilde{x}_{i+1}$ is equal to zero while $p^{k_{i-1}}a_{i}$ has order $p^{k_{i}-k_{i-1}}>p^{k_{i-1}}$. So for any $j\in\mathbb{N}$ the order of $\widetilde{x}_{1}$ is not less than $p^{k_{j}}$. Now note that
$k_{j}>2k_{j-1}>\ldots>2^{j-1}k_{1}\geqslant 2^{j-1}.$
It means that the order of $\widetilde{x}_{1}$ is not less than $p^{2^{j-1}}$ for every $j$. It is possible only if $\widetilde{x}_{1}$ has infinite order, but $A$ has no elements of infinite order, so the assumption of having a solution of the system in $A$ is false.
As a result, we have a unimodular system of equations over $A$ which has no solutions in $A$, as needed. ∎
Now let us move from $p$-groups to the other groups of unbounded period.
Lemma 2.
Suppose that $A$ is a reduced periodic abelian group of unbounded period. Then there is a unimodular system of equations over $A$ which has no solutions in $A$.
Proof.
The periodic abelian group $A$ is a direct sum of its $p$-components:
$A=\bigoplus_{i\in I}A_{p_{i}}.$
First consider the case where $I$ is finite. Then some of $A_{p}$ has unbounded period (otherwise the period of $A$ is bounded). If any unimodular system of equations over $A$ has a solution in $A$, then this is true for any quotient group of $A$. However, Proposition 1 shows that for $A_{p}$ this fact is not true. So, not all unimodular systems of equations over $A$ are solvable in $A$. One can even give an example of a system which has no solutions in $A$: if the projection of the system onto $A_{p}$ is the system from the proof of Proposition 1, then the system itself is unsolvable in $A$.
Now suppose that $I$ is infinite (in this case we can assume $I=\mathbb{N}$, although the numbers $p_{i}$ do not have to range over all prime numbers). In each $A_{p_{i}}$ pick an element $a_{i}\in A_{p_{i}}\setminus p_{i}A_{p_{i}}$, which exists since the abelian $p_{i}$-group $A_{p_{i}}$ is not divisible. Consider the following system of equations:
$\{x+p_{i}y_{i}=a_{i}\mid i\in\mathbb{N}\}.$
It has the following matrix of exponent sums:
$\begin{pmatrix}1&p_{1}&0&0&\dots\\ 1&0&p_{2}&0&\dots\\ 1&0&0&p_{3}&\dots\\ \dots&\dots&\dots&\dots&\dots\\ \end{pmatrix}$
Clearly, the system is $p_{i}$-nonsingular for every $p_{i}$ and $p$-nonsingular for the remaining prime numbers $p$, so this system is unimodular.
Suppose that this system has a solution $\{\widetilde{x},\widetilde{y}_{1},\widetilde{y}_{2},\ldots\}$ in $A$. It means that this solution lies in the Cartesian product $\prod\limits_{i\in\mathbb{N}}A_{p_{i}}$ of $p$-components. Look at the component $A_{p_{i}}$ and the equation $\widetilde{x}+p_{i}\widetilde{y}_{i}=a_{i}$. As $a_{i}\notin p_{i}A_{p_{i}}$, $p_{i}$-component of $\widetilde{x}$ is not 00 (otherwise $a_{i}$ is equal to $p_{i}\widehat{y}_{i}\in p_{i}A_{p_{i}}$, where $\widehat{y}_{i}$ is the $p_{i}$-component of $\widetilde{y}_{i}$). So we get that every component of $\widetilde{x}\in\prod\limits_{i\in\mathbb{N}}A_{p_{i}}$ is not equal to 00. As there is an infinite number of such components in $\prod\limits_{i\in\mathbb{N}}A_{p_{i}}$, $\widetilde{x}$ does not lie in the direct sum of the components, which means $\widetilde{x}\notin A$. We get a contradiction, so the assumption of finding a solution of the system in $A$ is false. ∎
2.2 Positive case
Now we look at groups which contain solutions to all unimodular systems of equations over themselves.
Proposition 2.
Any $p$-nonsingular system of equations over the abelian group $A$ of prime period $p$ has a solution in $A$ itself.
Proof.
Consider a $p$-nonsingular system $\{m_{i}=a_{i}\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over $A$, where $m_{i}$ is an integer linear combination of the variables $\{x_{j}\}$ (so $m_{i}$ is an element of the free $\mathbb{Z}$-module $\sum\limits_{j\in J}\mathbb{Z}\cdot x_{j}$), and $a_{i}$ is an element of $A$.
Note that the abelian group $A$ of period $p$ is a vector space over $\mathbb{Z}_{p}$ (and a $\mathbb{Z}$-module as well), whereas a solution of the system $\{m_{i}=a_{i}\}$ in $A$ is a homomorphism of $\mathbb{Z}$-modules $f\colon\sum\limits_{j\in J}\mathbb{Z}\cdot x_{j}\to A$ such that $f(m_{i})=a_{i}$.
Any such homomorphism passes through the $\mathbb{Z}$-module homomorphism $r_{p}\colon\sum\limits_{j\in J}\mathbb{Z}\cdot x_{j}\to\sum\limits_{j\in J}\mathbb{Z}_{p}\cdot x_{j}$, under which integer coefficients are taken modulo $p$. Therefore, we need to find a linear map $\widetilde{f}\colon\sum\mathbb{Z}_{p}\cdot x_{j}\to A$ such that $\widetilde{f}(\widetilde{m}_{j})=a_{j}$, where $\widetilde{m}_{j}=r_{p}(m_{j})$.
Now we construct the mapping $\widetilde{f}$. Note that vectors $\{\widetilde{m}_{i}\}$ are linearly independent over $\mathbb{Z}_{p}$ by the hypotheses of the proposition (as $m_{i}$ are the exponent sums rows of variables $\{x_{j}\}$ in $w_{i}$, and these rows are linearly independent modulo $p$), that is why there is a set $\{v_{k}\}_{k\in K}$ of vectors such that $\{\widetilde{m}_{i},v_{k}\}$ is a basis of $\sum\mathbb{Z}_{p}\cdot x_{j}$. Now define $\widetilde{f}$ on $\{\widetilde{m}_{i}\}\cup\{v_{k}\}$ so that $\widetilde{f}(\widetilde{m}_{i})=a_{i}$ and $\widetilde{f}(v_{k})$ are arbitrary (e.g. $\widetilde{f}(v_{k})=0$). We can now define $\widetilde{f}$ on $\sum\mathbb{Z}_{p}\cdot x_{j}$ as a linear map, and this is the desired linear map.
So, we get a solution of $\{m_{i}=a_{i}\}$ in $A$. ∎
Proposition 3.
Any $p$-nonsingular system of equations over an abelian $p$-group $A$ of bounded period is solvable in $A$ itself.
Proof.
Let the period of $A$ be $p^{n}$. Use induction on $n$.
The base case ($n=1$) is Proposition 2.
Now show that if the statement holds for $n=l$, then it holds for $n=l+1$. Consider a $p$-nonsingular system $\{m_{i}=a_{i}\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over $A$.
Look at the quotient group $A/pA$. Take the induced system $\{m_{i}=\overline{a}_{i}\}_{i\in I}$ of equations over $A/pA$, where $\overline{a}_{i}$ is the image of $a_{i}$ under taking the quotient $A\to A/pA$. It is a $p$-nonsingular system of equations over the abelian group $A/pA$ of period $p$, therefore by Proposition 2 the system has a solution $\{\bar{c}_{j}\}_{j\in J}$ in $A/pA$.
For each $\bar{c}_{j}\in A/pA$ pick an element $c_{j}\in A$ of its preimage. Then after the substitution $x_{j}\mapsto x_{j}+c_{j}$ the system $\{m_{i}=a_{i}\}$ transforms into $\{m_{i}=b_{i}\}$, where $b_{i}\in pA$, since the induced system $\{m_{i}=\overline{a}_{i}\}$ over $A/pA$ transforms into $\{m_{i}=1\}$.
The system $\{m_{i}=b_{i}\}$ can be seen as a system of equations over $pA$, which is an abelian group of period $p^{l}$. This system has a solution in $pA$ by induction hypothesis. This solution is also a solution of the system in $A\supset pA$, as needed. ∎
Now we can move on to arbitrary groups of bounded period.
Lemma 3.
Suppose that $A$ is an abelian group of bounded period. Then every unimodular system of equations over $A$ is solvable in $A$ itself.
Proof.
Suppose that $p_{1},p_{2},\ldots,p_{k}$ are all the prime divisors of the period of $A$. Then the following decomposition holds:
$A=A_{p_{1}}\oplus A_{p_{2}}\oplus\ldots\oplus A_{p_{k}},$
where $A_{p_{i}}$ is the $p_{i}$-component of $A$ (which has bounded period $p_{i}^{n_{i}}$). For the components $A_{p_{i}}$ the statement of the lemma follows from Proposition 3, therefore it is also true for their direct sum. ∎
Remark.
In fact, the system does not need to be precisely unimodular as long as it is $\{p_{1},\ldots,p_{k}\}$-nonsingular (since it suffices for system to be $p_{i}$-nonsingular for each prime $p_{i}$ which divides the period of the group).
Remark.
In particular, a unimodular (not necessarily finite) system of equations over a finite abelian group is always solvable in that group itself. This is also true for $\{p_{1},p_{2},\ldots,p_{k}\}$-nonsingular systems, where $p_{1},\ldots,p_{k}$ are all the prime divisors of the order of the abelian group.
2.3 The final criterion
Nowe we can prove the criterion formulated in the introduction.
Proof of Theorem 1.
Decompose $A$ into a direct sum of the divisible abelian group and the reduced one. The divisible abelian group has a solution of every nonsingular system of equations over itself (by Lemma 1). Therefore, nonsingular (particularly, unimodular) systems of equations have a solution in $A$ if and only if the same is true for its reduced part. Hence we can assume $A$ to be reduced.
By Lemma 2, if any unimodular system of equations over $A$ is solvable in $A$ itself, then $A$ must have bounded period.
And if $A$ has bounded period, then any unimodular system of equations over $A$ has a solution in $A$ by Lemma 3. ∎
Remark.
An analogous criterion can be formulated for abelian $\pi$-groups (a $\pi$-group is a group, elements of which have finite orders whose every prime divisor lies in $\pi$):
Suppose that $A$ is an abelian $\pi$-group, where $\pi$ is a set of some prime numbers. Then every $\pi$-nonsingular system of equations over $A$ is solvable in $A$ if and only if the reduced part of $A$ has bounded period.
3 Nilpotent groups
Now we switch back to the multiplicative notation. The obtained results can be used to research solvability of systems of equations in nilpotent groups.
3.1 Periodic nilpotent groups
Proof of Theorem 2.
Denote the nilpotency class of $G$ by $s$. We argue by induction on $s$.
The base case ($s=1$) follows from Theorem 1.
Now we show that if the statement is true for $s=l$, then it holds for $s=l+1$ as well. Consider a unimodular system $\{w_{i}=1\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over $G$.
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a unimodular system of equations over $G/Z(G)$ (which is a nilpotent group of class $l$ and of bounded period), therefore the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that the words $w_{i}$ initially have this property (that is, the products of their coefficients lie in $Z(G)$).
Search the solution $\{\widetilde{x}_{j}\}$ among the elements of $Z(G)$. Finding a solution of $\{w_{i}=1\}$, in which all $\widetilde{x}_{j}$ are from $Z(G)$, is the same as finding a solution of $\{v_{i}b_{i}=1\}_{i\in I}$ in $Z(G)$, where for each $i\in I$ the word $v_{i}$ is $w_{i}$ with coefficients removed. The latter system can be seen as a system of equations over $Z(G)$, thus (since the system is unimodular and $Z(G)$ has bounded period) it has a solution in $Z(G)$ by Theorem 1. This solution is also a solution of $\{w_{i}=1\}$, as needed. ∎
Remark.
Suppose that all the prime divisors of the period of the nilpotent group $G$ are $p_{1},p_{2},\ldots,p_{k}$. Then any $\{p_{1},p_{2},\ldots,p_{k}\}$-nonsingular system of equations over $G$ is solvable in $G$ as well.
For abelian groups it follows from the fact that the periods of primary components in the corresponding decomposition of such group can only be powers of prime numbers from $\{p_{1},p_{2},\ldots,p_{n}\}$.
For nilpotent groups we use the same argument as in the proof above.
In particular, an infinite-system analogue of Shmel’kin’s theorem (more precisely, of its existence part) holds for finite $p$-groups.
For finite nilpotent groups we can prove solvability of infinite unimodular systems of equations in another way:
Alternative proof.
Let $V$ be a unimodular system of equations over a finite nilpotent group $G$. For any finite subsystem $W\subset V$ consider the copy $G_{W}$ of $G$, which, by Shmel’kin’s theorem, has a solution $\{x_{W,j}\}_{j\in J}$ of $W$.
Now look at the set $M$ of all finite subsystems of $V$. Consider the following family of sets: by $S_{W}$ denote a set of all such finite systems $\widetilde{W}\subset V$ that $W\subset\widetilde{W}$. Note that $\{S_{W}\}$ has the finite intersection property. Indeed, the finite intersection $S_{W_{1}}\cap\ldots\cap S_{W_{n}}$ contains non-empty set $S_{W_{1}\cup\ldots\cup W_{n}}$ as a subset. Therefore, we can complete $\{S_{W}\}$ to an ultrafilter $\mathcal{U}$.
Now consider the ultraproduct $\left(\prod\limits_{W\in M}G_{W}\right)\Bigl{/}\mathcal{U}$ (in other words, the ultrapower of $G$). Embed $G$ into this ultrapower diagonally. The set $\{\widetilde{x}_{j}\}_{j\in J}$, where $\widetilde{x}_{j}$ is the element of the ultraproduct whose $G_{W}$-component is $x_{W,j}$ for each $W\in M$, forms a solution of $V$. Indeed, denote by $\widetilde{w}_{i}$ the word $w_{i}$ after the substitution of variables by the elements $\{\widetilde{x}_{j}\}$ of the ultraproduct. Then the set of indices of coordinates of $\widetilde{w}_{i}$, which are equal to $1$, contains $S_{\{w_{i}=1\}}$ as a subset, and so belongs the the ultrafilter. Hence $\widetilde{w}_{i}=1$ in the ultraproduct itself, and this holds for each $i\in I$.
But this ultraproduct is identical to $G$, as the order of an ultrapower of a finite group $G$ is not greater than $|G|$ [22, chapter IV, §8, item 8.5]. So the infinite unimodular system $V$ has a solution in $G$ itself. ∎
3.2 Divisible nilpotent groups
Before we prove the result concerning divisible nilpotent groups, let us prove the following lemma.
Lemma 4.
Suppose that $G$ is a divisible nilpotent group of class $s$. Then for every $i\in\{1,\ldots,s\}$ the term $Z_{i}(G)$ of upper central series of $G$ is divisible as well.
Proof.
Let us adopt the following convention: the commutator $[g,h]$ denotes $g^{-1}h^{-1}gh$.
Let $z$ be an element of $Z_{i}(G)$ and $n$ be a natural number.
Since $G$ is divisible, there is such an element $w\in G$ that $w^{n}=z$ Now we prove that $w\in Z_{i}(G)$. Let $g\in G$. There is an element $h\in G$ such that $h^{n}=g$.
Consider the commutator $[w,g]=[w,h^{n}]$. Assume that it is already shown that $w\in Z_{j}(G)$. This is true at least for $j=s$, as $Z_{s}(G)=G$. Using the commutator indentities
$[a,bc]=[a,c][a,b][[a,b],c];\quad[ab,c]=[a,c][[a,c],b][b,c],$
we get that
$[w,h^{n}]=[w,h^{n-1}][w,h][[w,h],h^{n-1}].$
Now notice that as $w\in Z_{j}(G)$, $[w,h_{1}]\in Z_{j-1}(G)$ and $[[w,h_{1}],h_{2}]\in Z_{j-2}(G)$ for all $h_{1},h_{2}\in G$. This means that $[w,h^{n}]=[w,h^{n-1}][w,h]$ modulo $Z_{j-2}(G)$. Dealing with $[w,h^{n-1}]$ in the same way and so on, we receive that $[w,h^{n}]=[w,h]^{n}$ modulo $Z_{j-2}(G)$.
Likewise we get that $[w^{n},h]=[w,h]^{n}$ modulo $Z_{j-2}(G)$. So, $[w,g]=[w,h^{n}]=[w,h]^{n}=[w^{n},h]=[z,h]$ modulo $Z_{j-2}(G)$. As $z\in Z_{i}(G)$, $[z,h]\in Z_{i-1}(G)$. Hence, $[w,g]\in Z_{i-1}(G)Z_{j-2}(G)$. If $j>i$, then $[w,g]\in Z_{j-2}(G)$ for any $g\in G$. So the element $w$ itself lies in $Z_{j-1}(G)$ (since it commutes modulo $Z_{j-2}(G)$ with any element of $G$).
As a result, we proved that if $w\in Z_{j}(G)$ and $j>i$, then $w\in Z_{j-1}(G)$. It follows that
$w\in Z_{s-1}(G),w\in Z_{s-2}(G),\ldots,w\in Z_{i}(G)$, hence the statement of the lemma. ∎
Remark.
For instance, the center of a divisible nilpotent group is divisible as well.
Remark.
In fact, we even prove that all terms of the upper central series of a divisible nilpotent group $G$ are isolated subgroups (i.e. any $n$th root, where $n$ is arbitrary, of any element of $Z_{i}(G)$ lies in $Z_{i}(G)$ as well).
Remark.
Notice that for $n$th root extraction in $Z_{i}(G)$ we needed only $n$th root extraction in $G$. Therefore, the analogous fact for $\pi$-divisibility (particularly, for $p$-divisibility) also holds:
Suppose that $G$ is a $\pi$-divisible nilpotent group. Then each of its upper central ceries terms is $\pi$-divisible as well.
Proof of Theorem 3.
Denote the nilpotency class of $G$ by $s$. We argue by induction on $s$.
The base case ($s=1$) follows from Lemma 1.
Now we show that if the statement is true for $s=l$, then it holds for $s=l+1$ as well. Consider a nonsingular system $\{w_{i}=1\}_{i\in I}$ of equations in variables $\{x_{j}\}_{j\in J}$ over $G$.
Take a look at the system $\{\overline{w}_{i}=1\}_{i\in I}$ over $G/Z(G)$, where for each $i\in I$ the word $\overline{w}_{i}$ is the word $w_{i}$ with coefficients changed to their images under taking the quotient $G\to G/Z(G)$. The system $\{\overline{w}_{i}=1\}$ is a nonsingular system of equations over $G/Z(G)$ (which is a divisible nilpotent group of class $l$), therefore the system is solvable in the quotient group $G/Z(G)$ itself by induction hypothesis.
This means that we can make a substitution of the variables in the system $\{w_{i}=1\}$ over $G$ such that for each $i\in I$ the product $b_{i}$ of coefficients of the word $w_{i}$ lies in $Z(G)$ (i.e. the image of $b_{i}$ under taking the quotient $G\to G/Z(G)$ is $1$). So we can assume that the words $w_{i}$ initially have this property (that is the products of their coefficients lie in $Z(G)$).
Search the solution $\{\widetilde{x}_{j}\}$ among the elements of $Z(G)$. Finding a solution of $\{w_{i}=1\}$, in which all $\widetilde{x}_{j}$ are from $Z(G)$, is the same as finding a solution of $\{v_{i}b_{i}=1\}_{i\in I}$ in $Z(G)$, where for each $i\in I$ the word $v_{i}$ is $w_{i}$ with coefficients removed. The latter system can be seen as a system of equations over $Z(G)$, thus (since the system is nonsingular and $Z(G)$ is divisible) it has a solution in $Z(G)$. This solution is also a solution of $\{w_{i}=1\}$, as needed. ∎
4 Open questions
Question 1.
Which torsion-free abelian groups contain solutions of every unimodular system of equations over itself? The group $\mathbb{Z}$ fails to have this property, as the following system of equations (in the additive notation) shows.
{,$\ 2y_{2}-y_{1}=0,\\ 2y_{1}-x=1,\\ x-3z_{1}=0,\\ z_{1}-3z_{2}=0,\\ \ldots\end{cases}$
Question 2.
The same question for mixed abelian groups.
Question 3.
Can a criterion similar to Theorem 1 be derived for nilpotent groups?